OpenStudy (anonymous):

enter in standard form 2x^2-3=-4x-1

OpenStudy (akashdeepdeb):

2x2 + 4x -2 = 0 Reduce it to x2 + 2x -1 + 0 Thus this is an identity = (x-1)^2 =0 Thus x = 1! :) I hope this helped! :)

OpenStudy (ivettef365):

that is 2x^2-3+4x+1 =0 2x^2+4x-2 = 0

OpenStudy (ivettef365):

Standard Form of a quadratic equation Ax^2 + Bx + C = 0

OpenStudy (akashdeepdeb):

Yeah ! :)

OpenStudy (ybarrap):

$$\tt 2x^2-3=-4x-1\\ 2x^2+4x-2=0$$